OpenStudy (anonymous):

A ball is dropped from a building 96 ft tall. The height of the ball after t seconds is measured by the function h(t) = −16t^2 + 96, where h(t) is measured in feet and t is time in seconds. How long will it take the ball to fall to 32 feet?

OpenStudy (turingtest):

h(t)=32=-16t^2+96 16t^2-64=0 t^2-4=0 t=2 sec

OpenStudy (phi):

double post? replace h(t) with 32 feet 32= -16t^2 +96 or 16t^2= 96-32 16t^2= 64 t^2= 64/16 can you finish?

OpenStudy (turingtest):

hey @phi, you think they mean h(t)=96-32?

OpenStudy (phi):

@turing: you did it right.

OpenStudy (turingtest):

I thought maybe they mean if it falls 32 feet it is at height 96-32=64 64=-16t^2+96 16t^2-32=0 t=sqrt(2)

OpenStudy (phi):

How long will it take the ball to fall *to* 32 feet which I interpret as h(t)= 32. h(0)= 96, which is where it starts.